3. Consider an editor that compares a typed in word say of length n and changes it to the nearest word in the dictionary, which can be of size m. For example, if you mistyped ”floyre”, it can change the word to ”flower” or ”flow” or ”floe” or ”floor”. To judge the best word, the spell checker computes the minimum number of changes between the typed word and the words in its dictionary. The changes can be as follows; (i) inserting a character, (ii) deleting a character or (iii) changing a character. All the change operations have equal weightage.

(a) Describe an algorithm to find the minimum number of changes between two given words. This is a well-known problem in dynamic programming, so you can take help of other resources. Just explain the process in your own words (20)

(b) Using this method show step by step how to determine to which word” sdimnas” should be changed. The candidates are (i)dynast, (ii) summer, (iii)dismal, (iv)dimmer, (v)sadden. (3*5=15) (please show the stapes too)

(c) Suggest two other criteria to select the most appropriate word, excluding the minimum number of changes. Gives reasons for why you selected the criteria (5)

4. Huffman Coding (40) The Latin alphabet is used in many European languages such as English, Spanish, French, German, etc. Show that the Huffman coding depends on the language used as well as the content. To do so apply Huffman coding on texts from different languages and different subjects, as follows;

Compress, using Huffman coding, the first 2000 words from Chapter 1 of the following novels in their original language; (Moby Dick; Don Quixote, Three Musketeers). (10)

Translate the exact 2000 words into English, Spanish, and French (whichever of the two is not the original language), using an online translator.

Compress these using Huffman code as well (10) Create a histogram comparing the length of the code per letter for (i) the three novels in the same languages and (ii) the same novel in different languages. Discuss your results
Answer #1

Answer 3:

We can solve this problem using dynamic programming. We can observe a pattern after solving some examples by hand that for writing each character we have two choices either get it by inserting or get it by copying, whichever takes less time. Now writing relation accordingly,
Let dp[i] be the optimal time to write i characters on screen then,

If i is even then,
   dp[i] = min((dp[i-1] + insert_time), 
               (dp[i/2] + copy_time))
Else (If i is odd)
   dp[i] = min(dp[i-1] + insert_time),
              (dp[(i+1)/2] + copy_time + removal_time) 

In the case of odd, removal time is added because when (i+1)/2 characters will be copied one extra character will be on the screen which needs to be removed. Total time complexity of solution will be O(N) and auxiliary space needed will be O(N).

whole code:

def minTimeForWritingChars(N, insrt,

                           remov, cpy):

    if N == 0:

        return 0

    if N == 1:

        return insrt

    dp = [0] * (N + 1)

    for i in range(1, N + 1):

        if i % 2 == 0:

            dp[i] = min(dp[i - 1] + insrt,

                        dp[i // 2] + cpy)

        else:

            dp[i] = min(dp[i - 1] + insrt,

                        dp[(i + 1) // 2] +

                        cpy + remov)

  

    return dp[N]